Hii i need help on this problem. And the question says write each expression as a simple fraction... -5 to the -3 power
Answer by stanbon(75887)

-5 to the -3 power
There are two ways to interprt your problem, as posted>
#1 if you meant -5^(-3) the
meaning is -(5^(-3))=-(1/5)^(3)=-1/125
#2 if you meant (-5)^(-3) the
meaning is (-1/5)^3= -1/125
The results are the same but you see the problem
could be misinterpreted.
Cheers,
Stan H.